OpenStudy (anonymous):

13. Find each measure listed below. Round all answers to the nearest tenth. (7 points) 10, 11, 4, 7, 12, 11, 16, 6, 9, 15 a. Mean: ___________ d. First Quartile: ___________ b. Median: ___________ e. Third Quartile: ___________ c. Mode: ___________ f. Standard Deviation: ___________ g. Variance: ___________

OpenStudy (anonymous):

The mode is the repeated value of 11. The others will require4 calculation.

OpenStudy (texaschic101):

4,6,7,9,10,11,11,12,15,16 mean = (4 + 6 + 9 + 11(2) + 12 + 15 + 16)/10 = 84/10 = 8.4 median (middle number ) aka Q2 = (10 + 11)/2 = 21/2 = 10.5 mode = 11 Q1 (first quartile) = (6 + 7)/2 = 13/2 = 6.5 Q3 (third quartile) = (12 + 15)/2 = 27/2 = 13.5 as for variance ans standard deviation... http://www.mathsisfun.com/data/standard-deviation.html

OpenStudy (anonymous):

okay thank you !(: